OpenStudy (anonymous):

If the points (1,3) (-3,-4) and (k, 3k-5) are collinear then what is/what are the value(s) of k?

OpenStudy (anonymous):

First find the slope.

OpenStudy (anonymous):

Then find the intercept.

OpenStudy (anonymous):

Then post your equation.

OpenStudy (anonymous):

y=(7/4)x+1.25

OpenStudy (anonymous):

I already found one value, k=5 but I was wondering if there were any other values since the question was stated the way it was

OpenStudy (anonymous):

You are right. k = 5. And all points on (x, (7/4)x+1.25) are on the line. if we include (k, 3k-5) as possible points, we have a system of equations in two variables. if k = x, then 3k-5 should = (7/4)x+1.25 sub in k for x on the right hand side 3k-5 = (7/4)k + 1.25 now we have one equation with one variable. That means there is only one value for k k=5

OpenStudy (anonymous):

I totally agree, thank you so much(:

OpenStudy (anonymous):

Happy to help :-) Study hard!